# RM-R1: Reward Modeling as Reasoning

### v2

@@ Appendix F Full Experiment Result
Cold-start RL evaluation <type>Chat</type> <rubric> 1. Accuracy of information (40 %): How correctly does the response describe the symptoms of sickle-cell disease? 2. Comprehensiveness (30 %): How thoroughly does the response cover the range of symptoms associated with sickle-cell disease? 3. Clarity and organization (20 %): How well-structured and easy to understand is the information presented? 4. Helpfulness for the user (10 %): How useful is the information for someone seeking to understand sickle-cell disease symptoms? <justify> These criteria focus on evaluating how well each response addresses the user’s question about sickle-cell disease symptoms. Accuracy is weighted highest because providing correct information about medical conditions is crucial. Comprehensiveness is important, as sickle-cell disease has multiple symptoms that vary in severity and presentation. Clarity and organization affect how easily the user can understand the information. Helpfulness considers how practical and actionable the information is for the user. </justify> </rubric> <eval> Comparing the two responses: Accuracy of information: <quote_A>1. Painful red or yellow skin lesions 2. Frequent infections 3. Fatigue and weakness 4. Anemia 5. Slower growth and development 6. Jaundice 7. Swelling of hands and feet 8. Vision loss 9. Tiredness and weakness after physical activity 10. Breathlessness 11. Fever, chills, and night sweats.</quote_A> <quote_B>1. Fatigue: People with sickle-cell disease often experience fatigue and weakness, even when they have not been physically active. 2. Pain: Sickle-cell disease can cause pain in the body, particularly in the arms, legs, back, and abdomen. Pain can be mild or severe and can last for days or weeks. 3. Fever: People with sickle-cell disease may be prone to infections and may experience frequent fevers. 4. Anemia: Sickle-cell disease can cause anemia, a condition in which the body does not have enough red blood cells to carry oxygen throughout the body. 5. Delayed growth and development: Children with sickle-cell disease may be smaller than their peers due to delayed growth and development. 6. Infections: People with sickle-cell disease are at increased risk for infections, particularly in the bloodstream and urinary tract. 7. Pigmentation: Sickle-cell disease can cause dark areas of pigmentation in the body, particularly on the hands and feet.</quote_B> Chatbot A provides a list of symptoms without explanation, but some listed items (e.g. “painful red or yellow skin lesions,” “vision loss”) are not typical of sickle-cell disease. Chatbot B offers more accurate information, explaining each symptom’s nature and severity. Comprehensiveness: Chatbot A lists 11 symptoms, several of which are inaccurate or uncommon. Chatbot B lists 7 symptoms that are more representative and supplies contextual detail about variability and medical advice. Clarity and organization: Chatbot A’s plain numbered list lacks context, making severity unclear. Chatbot B structures the list with explanations, improving readability and understanding. Helpfulness: Chatbot A is less helpful due to inaccuracies and missing context. Chatbot B is more useful, providing accurate descriptions, context, and guidance. </eval> <answer>[[B]]</answer>

### v3

@@ Appendix F Full Experiment Result
Table 9: Complete questions and chatbot responses.
Question: symptoms of sickle-cell disease Chatbot A: 1. Painful red or yellow skin lesions 2. Frequent infections 3. Fatigue and weakness 4. Anemia 5. Slower growth and development 6. Jaundice 7. Swelling of hands and feet 8. Vision loss 9. Tiredness and weakness after physical activity 10. Breathlessness 11. Fever, chills, and night sweats. Chatbot B Sickle cell disease is a genetic disorder that affects the shape of red blood cells. It is most common in African Americans, but can also affect people from other ethnic groups. The symptoms of sickle cell disease can vary widely from person to person, but some common symptoms include: 1. Fatigue: People with sickle cell disease often experience fatigue and weakness, even when they have not been physically active. 2. Pain: Sickle cell disease can cause pain in the body, particularly in the arms, legs, back, and abdomen. Pain can be mild or severe and can last for days or weeks. 3. Fever: People with sickle cell disease may be prone to infections and may experience frequent fevers. 4. Anemia: Sickle cell disease can cause anemia, a condition in which the body does not have enough red blood cells to carry oxygen throughout the body. 5. Delayed growth and development: Children with sickle cell disease may be smaller than their peers due to delayed growth and development. 6. Infections: People with sickle cell disease are at increased risk for infections, particularly in the bloodstream and urinary tract. 7. Pigmentation: Sickle cell disease can cause dark areas of pigmentation in the body, particularly on the hands and feet. It’s important to note that not everyone with sickle cell disease will experience all of these symptoms, and some people may not experience any symptoms at all. Therefore, it’s important to consult with a healthcare provider if you suspect you or someone you know may have sickle cell disease.
@@
Warm-start RL (RM-R1) evaluation <type>Chat</type> <rubric> 1. Accuracy of information (40 %): How correctly does the response describe the symptoms of sickle-cell disease? 2. Comprehensiveness (30 %): How thoroughly does the response cover the range of symptoms associated with sickle-cell disease? 3. Clarity and organization (20 %): How well-structured and easy to understand is the information presented? 4. Helpfulness for the user (10 %): How useful is the information for someone seeking to understand sickle-cell disease symptoms? <justify> These criteria focus on evaluating how well each response addresses the user’s question about sickle-cell disease symptoms. Accuracy is weighted highest because providing correct information about medical conditions is crucial. Comprehensiveness is important, as sickle-cell disease has multiple symptoms that vary in severity and presentation. Clarity and organization affect how easily the user can understand the information. Helpfulness considers how practical and actionable the information is for the user. </justify> </rubric> <eval> Comparing the two responses: Accuracy of information: <quote_A>1. Painful red or yellow skin lesions 2. Frequent infections 3. Fatigue and weakness 4. Anemia 5. Slower growth and development 6. Jaundice 7. Swelling of hands and feet 8. Vision loss 9. Tiredness and weakness after physical activity 10. Breathlessness 11. Fever, chills, and night sweats.</quote_A> <quote_B>1. Fatigue: People with sickle-cell disease often experience fatigue and weakness, even when they have not been physically active. 2. Pain: Sickle-cell disease can cause pain in the body, particularly in the arms, legs, back, and abdomen. Pain can be mild or severe and can last for days or weeks. 3. Fever: People with sickle-cell disease may be prone to infections and may experience frequent fevers. 4. Anemia: Sickle-cell disease can cause anemia, a condition in which the body does not have enough red blood cells to carry oxygen throughout the body. 5. Delayed growth and development: Children with sickle-cell disease may be smaller than their peers due to delayed growth and development. 6. Infections: People with sickle-cell disease are at increased risk for infections, particularly in the bloodstream and urinary tract. 7. Pigmentation: Sickle-cell disease can cause dark areas of pigmentation in the body, particularly on the hands and feet.</quote_B> Chatbot A provides a list of symptoms without explanation, but some listed items (e.g. “painful red or yellow skin lesions,” “vision loss”) are not typical of sickle-cell disease. Chatbot B offers more accurate information, explaining each symptom’s nature and severity. Comprehensiveness: Chatbot A lists 11 symptoms, several of which are inaccurate or uncommon. Chatbot B lists 7 symptoms that are more representative and supplies contextual detail about variability and medical advice. Clarity and organization: Chatbot A’s plain numbered list lacks context, making severity unclear. Chatbot B structures the list with explanations, improving readability and understanding. Helpfulness: Chatbot A is less helpful due to inaccuracies and missing context. Chatbot B is more useful, providing accurate descriptions, context, and guidance. </eval> <answer>[[B]]</answer>